19. Joe bought a house for $200,000 with $50,000 down payment and the rest an 8% 20 year constant payment mortgage. If his payment is $1,280.27 per month, how many discount points the lender charged on the loan?

A. 0 pt

B. 1 pt

C. 2 pt

D. 3 pt

20. In the above question, if the lender also charges a 5% prepayment penalty if the loan is paid off before its maturity, how much does Joe owe to the bank if he decides to pay off the loan the same time he makes the 60th monthly payments?

      A.   $131,288

      B.   $140,666

      C.   $137,852

      D.   $133,968

21. In question 19 and 20, what is the effective interest rate of the mortgage when Joe pays off the mortgage at the end of 60th month?

      A. 8%

      B    8.55%

      C.   9.26%
Answer #1

19) Loan value for 8% 20 yr mortgage with monthly payment of 1280.27

Loan value $153,061.77 =-PV(8%/12,20*12,1280.27,,)

Total cost 200000.00

Downpayment 50000.00

Loan without discount points 150000.00

Discount points charged 2.04%

So answer is c) 2 pt

20) Value of loan after 60th installment 133968.21 =FV(8%/12,60,1280.27,-153061.77,)

5% penalty 6698.41

Total 140666.62

Correct answer is B) 140,666

21) Effective interest rate 9.26% =RATE(60,1280.27,-150000,140666,,)*12

Correct answer is C) 9.26%